# Is the Faber-Krahn inequality true for the Stokes operator?

### v1

@@ 1.3 Bibliographical references
Due to the wealth of information they provide on the interplay between the geometry and the analytic properties (i.e. regarding functions defined on the domain) of domains, spectral optimisation problems have become a tenet of applied mathematics. The basic question in spectral optimisation is the following: given a certain spectral quantities, defined through a differential operator on a domain, what is the domain minimising or maximising this spectral quantity? It would be pointless to try and give an exhaustive bibliographical account of the developments of the field, but let us highlight some of the key contributions. The seminal works of Faber [23] and Krahn [33], investigating the minimisation of the first eigenvalue of the (scalar) Dirichlet-Laplacian bolstered the development of new approaches blending geometric and analytic tools. Typical questions include the existence of optimisers, their geometry and their stability. Regarding the existence of optimisers, when no direct comparison principles can yield the explicit description of an optimiser, the first general theorem is due to Buttazzo & Dal Maso [10], when additional box constraints are enforced on the set of admissible domains. Bucur [3] developed a framework designed to handle the unbounded case. This was done using the concentration compactness-principle of Lions [37], and later used by Bucur & Varchon [6] and Bucur & Freitas [5] to derive existence results for the optimisation of eigenvalues of scalar operators. Let us also mention some more general results due, independently, to Bucur and Mazzoleni & Pratelli [4, 38]. We refer to [28, chapter 2] as a reference for existence problems. Regarding the geometry of optimisers, these are usually very delicate questions. While symmetrisation techniques-which can be used to derive direct comparison results-can be available for specific problems, there are no general tools that can be employed for generic spectral optimisation problems. Regarding rearrangements, we refer to the monograph [32]. Another possibility to obtain information regarding the optimal domains is to derive tractable optimality conditions; as the latter are often expressed in terms of an overdetermined boundary value problem having a solution on the optimal domain, results similar to the Serrin theorem [45] (see also [40]) yield the result. For an introduction to the geometric aspects of spectral optimisation we refer to [29].
@@
The literature devoted to shape optimisation problems for vectorial operators is scarce, but growing. Of particular importance to us in this regard are the different contributions to the study of isoperimetric inequalities for the ${\mathbf{curl}}$ operator, that is, the optimisation of the first positive eigenvalue of
$\begin{cases}{\mathbf{curl}}({\mathbf{u}})=\xi{\mathbf{u}}&\text{ in }{\Omega}\,,\\ \langle u,\nu\rangle=0&\text{ on }\partial{\Omega}.\end{cases}$
This problem is of paramount importance in the study of magnetic fields, and has a strong connection [14] to the Stokes eigenvalue problem with a tangential flow condition, rather than a Dirichlet condition. To the best of our knowledge, the main works on this ${\mathbf{curl}}$-isoperimetric problems can be found, one the one hand, in the works of Cantarella, DeTurck, Gluck and Teytel [11, 12] and, on the other hand, in the research of Enciso, Gerner & Peraltas-Sala [21, 22, 25]. Let us briefly review their finding: in [11, 12], the authors, using explicit computations of eigenfunctions, observe the optimality of the ball among all concentric spherical shells with a given volume constraint, but also conjecture that the minimiser of the ${\mathbf{curl}}$-isoperimetric problem is a “spheromak”, that is, a sphere where the south and north poles are glued together. In their paper, they also give optimality conditions for the optimiser, which lead to the same type of conclusion as Corollary 1. It is interesting to note that their analysis recovers parts of the semi-differentiability results of [15]. In [21, 22], another take on the problem is introduced, and the authors investigate possible symmetries of optimisers. In [22], Enciso & Peraltas-Sala investigate whether a solution of the ${\mathbf{curl}}$-isoperimetric problem can have axial symmetry; their conclusion is no, provided some regularity of the minimiser is assumed a priori. Their results hinge on a variational analysis (à la Hadamard). In [21] on the other hand, they seek optimisers in the class of convex domains, and they prove that one can not have “too regular” optimisers. Gerner, in [25], went much further in the fine characterisation of possible optimisers, both in the euclidean and riemannian settings . Although the adaptation of some of his results is not immediate, one of the necessary optimality conditions we derive is inspired by [25, Theorem 2]. Finally, let us mention the recent work of Lamberti & Zaccaron [35], in which, also using a semi-differential approach, the authors investigate the optimal shape of an electromagnetic cavity.
@@
for some function $\psi$. Taking the (2-dimensional) ${\mathbf{curl}}$ of the equation, it appears that $(\psi,\lambda_{1}({\Omega}))$ actually solves
@@
We refer to [27, Chapter 11] for more details, but let us underline the following aspects of the problem: first of all, the existence of an optimal domain remained open until the contribution of Ashbaugh & Bucur [1] and, later, of Stollenwerk [48, 47]. Second, this problem is related to a long-standing conjecture by Pólyà & Szegö that asserts that the ball is a solution of this variational problem. A very active line of research has focused on the computation of shape-derivatives for this problem and, more generally, for polyharmonic problems. Regarding the Hadamard type shape derivatives, we refer to [26]. For applications of this calculus to polyharmonic problems, let us point to the numerous works of Buoso & Lamberti [7, 8, 9]. To the best of our knowledge, the second-order derivative of the first eigenvalue at the ball was not known.
@@ 2 Proof of Theorem 1
Our proof is in two steps. We first apply the concentration-compactness principle in order to show that one might, in (8), further assume ${\Omega}\subset D$ for a fixed compact set $D\subset\mathbb{R}^{d}$. We then investigate the existence of an optimiser with this additional box constraint.
Step 1: Using the.
@@
$\int_{\mathbb{R}^{d}}w_{k}^{2}=1,\quad\int_{\mathbb{R}^{d}}|{\nabla}w_{k}|^{2}\leqslant\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{u}}_{k}\|^{2}.$ (9)
By the concentration-compactness principle of Lions [37] we know that one of the following occurs:
@@
    Concentration: there exists a function $w\in L^{2}(\mathbb{R}^{d};\mathbb{R}_{+})$ and a sequence $\{y_{k}\}_{k\in\mathbb{N}}\in(\mathbb{R}^{d})^{\mathbb{N}}$ such that $w_{k}(\cdot+y_{k})\xrightarrow[k\to\infty]{}w$ in $L^{2}(\mathbb{R}^{d})$.
@@
$\psi_{k,1}(x)=\begin{cases}1\text{ in }{\mathbb{B}}(y_{k},R_{k})\,,\\ 0\text{ in }{\mathbb{B}}(y_{k},R_{k}+2\eta_{k})^{c},\end{cases}\psi_{k,2}(x)=\begin{cases}1\text{ in }{\mathbb{B}}(y_{k},R_{k}^{\prime}-\eta_{k})^{c}\,,\\ 0\text{ in }{\mathbb{B}}(y_{k},R_{k}^{\prime}-2\eta_{k})\end{cases}$
@@
In particular,
@@
As $\phi$ is compactly supported, $\mathrm{supp}(\phi)\cap\{{\nabla}\cdot{\mathbf{v}}_{k,1}\neq 0\}=\emptyset$ for $k$ large enough. Indeed, should this not be the case, as ${\nabla}\cdot{\mathbf{v}}_{k,1}=0$ in ${\mathbb{B}}(y_{k},R_{k}+\eta_{k})$, it would follow that
$\mathrm{dist}(\mathrm{supp}(\phi),{\mathbb{B}}(y_{k},R_{k}))\geqslant\eta_{k}\xrightarrow[k\to+\infty]{}\infty.$
This contradicts the strong $L^{2}$ convergence of $\{{\mathbf{v}}_{k,1}\}_{k\in\mathbb{N}}$, thereby leading to a contradiction. Thus,
@@
and, for any $j\in\mathbb{N}$, we have a sequence $\{{\mathbf{v}}_{k,2}^{j}\}_{k\in\mathbb{N}\,,j\leqslant J_{k}}$ where $J_{k}\xrightarrow[k\to+\infty]{}\infty$ such that for any $j$, $\{{\mathbf{v}}_{k,2}^{j}\}_{k\in\mathbb{N}}$ is in a situation of concentration, and $\{\{{\mathbf{v}}_{k,2}^{J_{k}}\neq 0\}\}_{k\in\mathbb{N}}$ is a minimising sequence.
@@
$\lambda^{*}\geqslant\underset{k\to\infty}{\lim\inf}\frac{\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,1}\|^{2}+\sum_{j=1}^{J_{k}}\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,2}^{j}\|^{2}}{\sum_{j=1}^{J_{k}}\alpha_{j}}>\lambda^{*}$
by convexity since we have, by construction,
$\underset{k\to\infty}{\lim\inf}\frac{\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,1}\|^{2}}{\alpha_{1}}>\lambda^{*}.$
This is a contradiction, and the result follows.
We deduce that we have concentration of the sequence. Upon examination of the truncation procedure, one can see that the condition $R_{k}\underset{k\to\infty}{\rightarrow}+\infty$ can be weakened, and that $\{R_{k}\}_{k\in\mathbb{N}}$ might be chosen to be constant, as is the case in the classical statement of the concentration-compactness lemma [37]. This allows to introduce a compact set $D\subset\mathbb{R}^{d}$, and to only consider sequences $\{\Omega_{k}\}_{k\in\mathbb{N}}\subset D$, that is, we can now simply deal with an optimisation problem with box constraints.
Step 2: Existence under box-constraints.
We consider a fixed compact domain $D\subset\mathbb{R}^{d}$. Let us define
$\widehat{\mathscr{O}}_{V_{0}}=$Ω quasi-open included in $\left\{\Omega\textrm{ quasi-open included in }D\,,|\Omega|\leqslant V_{0}\right\}$
and let us introduce, for $\Omega\in\widehat{\mathcal{O}}_{V}$, the quantity $\widehat{\lambda_{1}}(V)$ given by
$\widehat{\lambda_{1}}(V_{0})=\inf_{\begin{subarray}{c}\mathbf{u}\in W^{1,2}_{0}(D)^{3}\\ \int_{D}\mathbf{u}\cdot\nabla v=0\ \forall v\in H^{1}(D)\\ |\Omega_{\mathbf{u}}|\leqslant V_{0}\\ \int_{\Omega}\|{\mathbf{u}}\|^{2}=1\end{subarray}}{\int_{D}\|\nabla\mathbf{u}\|^{2}}$ (10)
where $\Omega_{\mathbf{u}}$ denotes the quasi open set $\{\mathbf{u}\neq 0\}$.
Let $\{{\mathbf{u}}_{k}\}_{k\in\mathbb{N}}$ be minimising for Problem (10). For any $k\in\mathbb{N}$, let $\Omega_{k}:=\{{\mathbf{u}}_{k}\neq 0\}$. In view of the Poincaré inequality [30, Lemma 4.5.3], $\{{\mathbf{u}}_{k}\}_{k\in\mathbb{N}}$ is bounded in $W^{1,2}_{0}(D)^{3}$. It follows that there exists $\mathbf{u}^{\star}\in W^{1,2}_{0}(D)^{3}$ such that $\{{\mathbf{u}}_{k}\}_{k\in\mathbb{N}}$ converges to $\mathbf{u}^{\star}$ up to a (not relabelled) subsequence, weakly in $W^{1,2}_{0}$ and strongly in $L^{2}$. Thus, $\{u_{k}\}_{k\in\mathbb{N}}$ converges almost everywhere to $\mathbf{u}^{\star}$, and therefore
$|\Omega_{\mathbf{u^{\star}}}|\leqslant\liminf_{k\to+\infty}|\Omega_{k}|\leqslant V_{0}.$
where $\Omega_{\mathbf{u^{\star}}}$ denotes the quasi-open set $\{\mathbf{u^{\star}}\neq 0\}$. We infer that $\Omega_{\mathbf{u^{\star}}}$ belongs to $\widehat{\mathscr{O}}_{V}$. Furthermore, since $\mathbf{u^{\star}}=0$ quasi-everywhere on $D\backslash\Omega_{\mathbf{u}^{\star}}$, by weak $H^{1}$-convergence, one has
by minimality, whence the equality of these quantities. By strong convergence in $L^{2}(D;\mathbb{R}^{3})$ and weak convergence in $W^{1,2}_{0}(D)^{3}$, there holds
$\int_{D}\|\mathbf{u}^{*}\|^{2}=1$
and, for any $v\in W^{1,2}(D)$,
Thus, ${\Omega}_{{\mathbf{u}}^{*}}$ is a solution of (10). The proof is now concluded.

### v2

@@ 1.3 Bibliographical references
Due to the wealth of information they provide on the interplay between the geometry and the analytic properties (i.e. regarding functions defined on the domain) of domains, spectral optimisation problems have become a tenet of applied mathematics. The basic question in spectral optimisation is the following: given a certain spectral quantity, defined through a differential operator on a domain, what is the domain minimising or maximising this spectral quantity? It would be pointless to try and give an exhaustive bibliographical account of the developments of the field, but let us highlight some of the key contributions. The seminal works of Faber [23] and Krahn [38], investigating the minimisation of the first eigenvalue of the (scalar) Dirichlet-Laplacian bolstered the development of new approaches blending geometric and analytic tools. Typical questions include the existence of optimisers, their geometry and their stability. Regarding the existence of optimisers, when no direct comparison principles can yield the explicit description of an optimiser, the first general theorem is due to Buttazzo & Dal Maso [10], when additional box constraints are enforced on the set of admissible domains. Bucur [3] developed a framework designed to handle the unbounded case. This was done using the concentration compactness-principle of Lions [44], and later used by Bucur & Varchon [6] and Bucur & Freitas [5] to derive existence results for the optimisation of eigenvalues of scalar operators. Let us also mention some more general results due, independently, to Bucur and Mazzoleni & Pratelli [4, 45]. We refer to [32, chapter 2] as a reference for existence problems. Regarding the geometry of optimisers, these are usually very delicate questions. While symmetrisation techniques-which can be used to derive direct comparison results-can be available for specific problems, there are no general tools that can be employed for generic spectral optimisation problems. Regarding rearrangements, we refer to the monograph [37]. Another possibility to obtain information regarding the optimal domains is to derive tractable optimality conditions; as the latter are often expressed in terms of an overdetermined boundary value problem having a solution on the optimal domain, results similar to the Serrin theorem [52] (see also [47]) yield the result. For an introduction to the geometric aspects of spectral optimisation we refer to [33].
@@
The literature devoted to shape optimisation problems for vectorial operators is scarce, but growing. Of particular importance to us in this regard are the different contributions to the study of isoperimetric inequalities for the ${\mathbf{curl}}$ operator in $\mathbb{R}^{3}$, that is, the optimisation of the first positive eigenvalue of
$\begin{cases}{\mathbf{curl}}({\mathbf{u}})=\xi{\mathbf{u}}&\text{ in }{\Omega}\,,\\ \langle{\mathbf{u}},\nu\rangle=0&\text{ on }\partial{\Omega}.\end{cases}$
This problem is of paramount importance in the study of magnetic fields, and has a strong connection [14] to the Stokes eigenvalue problem with a tangential flow condition, rather than a Dirichlet condition. To the best of our knowledge, the main works on this ${\mathbf{curl}}$-isoperimetric problems can be found, on the one hand, in the works of Cantarella, DeTurck, Gluck and Teytel [11, 12] and, on the other hand, in the research of Enciso, Gerner & Peralta-Salas [21, 22, 28]. Let us briefly review their finding: in [11, 12], the authors, using explicit computations of eigenfunctions, observe the optimality of the ball among all concentric spherical shells with a given volume constraint, but also conjecture that the minimiser of the ${\mathbf{curl}}$-isoperimetric problem is a “spheromak”, that is, a sphere where the south and north poles are glued together. In their paper, they also give optimality conditions for the optimiser, which lead to the same type of conclusion as Corollary 1. It is interesting to note that their analysis recovers parts of the semi-differentiability results of [15]. Before we describe the results of [21, 22, 28] observe that Enciso, Gerner & Peralta-Salas work with vector fields that are not only incompressible, but also orthogonal to harmonic (i.e. ${\mathbf{curl}}$ and ${\mathrm{div}}$ free) fields. In [21, 22], another take on the problem is introduced, and the authors investigate possible symmetries of optimisers. In [22], Enciso & Peralta-Salas investigate whether a solution of the ${\mathbf{curl}}$-isoperimetric problem can have axial symmetry; their conclusion is no, provided some regularity of the minimiser is assumed a priori. Their results hinge on a variational analysis (à la Hadamard). In [21] on the other hand, they seek optimisers in the class of convex domains, and they prove that one can not have “too regular” optimisers. Gerner, in [28], went much further in the fine characterisation of possible optimisers, both in the euclidean and riemannian settings . Although the adaptation of some of his results is not immediate, one of the necessary optimality conditions we derive is inspired by [28, Theorem 2]. Finally, let us mention the recent work of Lamberti & Zaccaron [40], in which, also using a semi-differential approach, the authors investigate the optimal shape of an electromagnetic cavity.
@@
for some function $\psi$. Plugging this expression into (2) we deduce that
$\begin{cases}-\partial_{xxy}\psi-\partial_{yyy}\psi-\partial_{x}p=\lambda_{1}({\Omega})\partial_{y}\psi&\text{ in }{\Omega}\,,\\ -\partial_{xxx}\psi-\partial_{yyx}\psi+\partial_{y}p=\lambda_{1}({\Omega})\partial_{x}\psi&\text{ in }{\Omega}.\end{cases}$
Differentiating the first equation with respect to $y$, the second with respect to $x$ and summing the two, we deduce that $-\Delta^{2}\psi=\lambda_{1}({\Omega})\Delta\psi$. Furthermore, as $\begin{pmatrix}-\partial_{y}\psi\\ \partial_{x}\psi\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix}$ on $\partial{\Omega}$ we deduce that $\psi$ is constant on $\partial{\Omega}$, and that $\partial_{\nu}\psi=0$. Up to adding a constant to $\psi$, we may take $\psi=0$ on $\partial{\Omega}$. Overall, $(\psi,\lambda_{1}({\Omega}))$ solves the fourth order equation
@@
We refer to [31, Chapter 11] for more details, but let us underline the following aspects of the problem: first of all, the existence of an optimal domain remained open until the contribution of Ashbaugh & Bucur [1] and, later, of Stollenwerk [55, 54]. Second, this problem is related to a long-standing conjecture by Pólyà & Szegö that asserts that the ball is a solution of this variational problem. A very active line of research has focused on the computation of shape-derivatives for this problem and, more generally, for polyharmonic problems. Regarding Hadamard type shape derivatives, we refer to [30]. For applications of this calculus to polyharmonic problems, let us point to the numerous works of Buoso & Lamberti [7, 8, 9]. To the best of our knowledge, the second-order derivative of the first eigenvalue at the ball was not known.
@@ 2 Proof of Theorem 1
$\int_{\mathbb{R}^{d}}w_{k}^{2}=\int_{\Omega}\|{\mathbf{u}}_{k}\|^{2}\color[rgb]{0,0,0}=1,\quad\int_{\mathbb{R}^{d}}|{\nabla}w_{k}|^{2}\leqslant\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{u}}_{k}\|^{2}.$ (9)
By the concentration-compactness principle of Lions [44] we know that, up to extracting a subsequence which we still denote $\{w_{k}\}_{k\in\mathbb{N}}$ with a slight abuse of notations, one of the following occurs:
@@
    Concentration: there exists a function $w\in L^{2}(\mathbb{R}^{d};\mathbb{R}_{+})$ and a sequence $\{y_{k}\}_{k\in\mathbb{N}}\in(\mathbb{R}^{d})^{\mathbb{N}}$ such that $w_{k}(\cdot+y_{k})\xrightarrow[k\to\infty]{}w$ in $L^{2}(\mathbb{R}^{d})$ and weakly in $W^{1,2}(\mathbb{R}^{d})$.
@@
$\psi_{k,1}(x)=\begin{cases}1\text{ in }{\mathbb{B}}(y_{k},R_{k}+\eta_{k}\color[rgb]{0,0,0})\,,\\ 0\text{ in }{\mathbb{B}}(y_{k},R_{k}+2\eta_{k})^{c},\end{cases}\psi_{k,2}(x)=\begin{cases}1\text{ in }{\mathbb{B}}(y_{k},R_{k}^{\prime}-\eta_{k})^{c}\,,\\ 0\text{ in }{\mathbb{B}}(y_{k},R_{k}^{\prime}-2\eta_{k})\end{cases}$
@@
Then, taking into account that $\|{\nabla}\psi_{k,1}\|_{L^{\infty}(\mathbb{R}^{d})}\,,\|{\nabla}\psi_{k,1}\|_{L^{\infty}(\mathbb{R}^{d})}\leqslant 1/\eta_{k}$, we can proceed exactly as in [44, Lemma III.1] to obtain
@@
As $\phi$ is compactly supported, $\mathrm{supp}(\phi)\cap\{{\nabla}\cdot{\mathbf{v}}_{k,1}\neq 0\}=\emptyset$ for $k$ large enough. Indeed, should this not be the case, as ${\nabla}\cdot{\mathbf{v}}_{k,1}={\nabla}\cdot{\mathbf{u}}_{k}=0$ in ${\mathbb{B}}(y_{k},R_{k}+\eta_{k})$, we conclude that there exists $x\in\mathrm{supp}(\phi)\cap{\mathbb{B}}(y_{k},R_{k}+\eta_{k})^{c}$. Letting $\mathrm{dist}(E,F)$ be the Hausdorff distance between two closed subset $E\,,F\subset\mathbb{R}^{d}$, this implies
$\mathrm{dist}\left(\mathrm{supp}(\phi),{\mathbb{B}}(y_{k},R_{k})\right)\geqslant\eta_{k}\xrightarrow[k\to+\infty]{}\infty.$
In particular, we deduce that
${\mathbf{v}}_{k,1}\underset{k\to\infty}{\rightarrow}0\text{ in }L^{2}_{\mathrm{loc}}(\mathbb{R}^{d}),$
which contradicts the strong $L^{2}$ convergence of $\{{\mathbf{v}}_{k,1}\}_{k\in\mathbb{N}}$, thereby leading to a contradiction. Thus,
@@
and we have a sequence $\{{\mathbf{v}}_{k,2}^{j}\}_{k\in\mathbb{N}\,,j\leqslant J_{k}}$ where $J_{k}\xrightarrow[k\to+\infty]{}\infty$ such that for any $j$, $\{{\mathbf{v}}_{k,2}^{j}\}_{k\in\mathbb{N}}$ is in a situation of concentration, and $\{\{{\mathbf{v}}_{k,2}^{J_{k}}\neq 0\}\}_{k\in\mathbb{N}}$ is a minimising sequence.
@@
$\lambda^{*}\geqslant\underset{k\to\infty}{\lim\inf}\frac{\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,1}\|^{2}+\sum_{j=1}^{J_{k}}\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,2}^{j}\|^{2}}{\int_{\mathbb{R}^{d}}\|{\mathbf{v}}_{k,1}\|^{2}+\sum_{j=1}^{J_{k}}\int_{\mathbb{R}^{d}}\|{\mathbf{v}}_{k,2}^{j}\|^{2}}.$
Let
$\alpha_{k,1}:=\int_{\mathbb{R}^{d}}\|{\mathbf{v}}_{k,1}^{j}\|^{2}\text{ which satisfies }\alpha_{k,1}\underset{k\to\infty}{\rightarrow}\alpha_{1}>0$
and
$\tilde{\alpha}_{k}:=\sum_{j=1}^{J_{k}}\int_{\mathbb{R}^{d}}\|{\mathbf{v}}_{k,2}^{j}\|^{2}\text{ which satisfies }\tilde{\alpha}_{k,1}\underset{k\to\infty}{\rightarrow}1-\alpha_{1}.$
Thus
$\lambda^{*}\geqslant\underset{k\to\infty}{\lim\inf}\frac{\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,1}\|^{2}}{\alpha_{k,1}+\tilde{\alpha}_{k}}+\underset{k\to\infty}{\lim\inf}\frac{\sum_{j=1}^{J_{k}}\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,2}^{j}\|^{2}}{\alpha_{k,1}+\tilde{\alpha}_{k}}.$
Now, since $\alpha_{k,1}+\tilde{\alpha}_{k}\underset{k\to\infty}{\rightarrow}1\,,\alpha_{k,1}\underset{k\to\infty}{\rightarrow}\alpha_{1}$, the fact that we have by construction
$\underset{k\to\infty}{\lim\inf}\frac{\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,1}\|^{2}}{\alpha_{1}}>\lambda^{*}$
implies
$\underset{k\to\infty}{\lim\inf}\frac{\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,1}\|^{2}}{\alpha_{k,1}+\tilde{\alpha}_{k}}>\alpha_{1}\lambda^{*}.$
Likewise,
$\underset{k\to\infty}{\lim\inf}\frac{\sum_{j=1}^{J_{k}}\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,2}^{j}\|^{2}}{\alpha_{k,1}+\tilde{\alpha}_{k}}\geqslant(1-\alpha_{1})\lambda^{*}.$
Consequently
$\lambda^{*}\geqslant\underset{k\to\infty}{\lim\inf}\frac{\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,1}\|^{2}+\sum_{j=1}^{J_{k}}\int_{\mathbb{R}^{d}}\|{\nabla}{\mathbf{v}}_{k,2}^{j}\|^{2}}{\int_{\mathbb{R}^{d}}\|{\mathbf{v}}_{k,1}\|^{2}+\sum_{j=1}^{J_{k}}\int_{\mathbb{R}^{d}}\|{\mathbf{v}}_{k,2}^{j}\|^{2}}>\lambda^{*},$
a contradiction. The result follows.
We deduce that we have concentration of the sequence, so that
${\mathbf{v}}_{k}\underset{k\to\infty}{\rightarrow}{\mathbf{v}}\text{ in $L^{2}(\mathbb{R}^{d})$}.$
Define
${\Omega}^{*}:=\{{\mathbf{v}}\neq 0\}.$
[…]
$\lambda_{1}({\Omega}^{*})\leqslant\frac{\int_{{\Omega}^{*}}\|{\nabla}{\mathbf{v}}\|^{2}}{\int_{\mathbb{R}^{d}}\|{\mathbf{v}}\|^{2}}\leqslant\underset{k\to\infty}{\lim\inf}\frac{\int_{{\Omega}^{*}}\|{\nabla}{\mathbf{v}}_{k}\|^{2}}{\int_{\mathbb{R}^{d}}\|{\mathbf{v}}_{k}\|^{2}}=\lambda^{*},$
where we used the weak $W^{1,2}(\mathbb{R}^{d};\mathbb{R}^{d})$ convergence of the sequence $\{{\mathbf{v}}_{k}\}_{k\in\mathbb{N}}$.
It follows that $\Omega^{*}$ solves  (8).